# Methodology, Software, Validation, Writing-Original Draft, Writing-Review and Editing, Visualization

### v1

@@ 4 Conclusion
In this work, we proposed a hybrid neural network model to hierarchically extract features on the functional brain networks for ASD identification. Extensive experiments demonstrate the effectiveness of the proposed model. Compared with other baseline methods, the proposed model achieve superior performance on widely used dataset ABIDE-I. The ablation experiments and the interpretability analysis further verify the rationality of the model architecture. The generalization performance was also explored preliminarily on the ADHD-200 dataset. In future studies, it is benifical to integrate neural mechanisms into the deep learning models to boost the ASD identification.

### v2

@@ 4 Conclusion
In this work, we proposed a hybrid neural network model to hierarchically extract features on the functional brain networks for ASD identification. Extensive experiments demonstrate the effectiveness of the proposed model. Compared with other baseline methods, the proposed model achieve superior performance on widely used dataset ABIDE-I. The ablation experiments and the interpretability analysis further verify the rationality of the model architecture. The generalization performance was also explored preliminarily on the ADHD-200 dataset and ABIDE-II dataset. In future studies, it is beneficial to integrate neural mechanisms into the deep learning models to boost the ASD identification.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No.62076209.